Related Rates: Problem Solving


Direction: Solve each word problem involving a related rates. Make a sketch of your work

if necessary, and then apply differentiation.


A light is hung 15 feet above a straight horizontal path. If a man 6 feet tall walking

away from the light at a rate of 5 feet per second, how fats is the man’s shadow

lengthening when the man is 8 feet from the base of the light source?


PLEASE ANSWER MY QUESTION QUICKLY!!

DEADLINE : 05/03/2022 11 : 00 PM








1
Expert's answer
2022-05-03T04:11:52-0400

x is the distance from the man to the pole, and y is the distance from the tip of the man's shadow to the pole. i assume the man and pole are standing straight up, which means the 2 triangles are similar.

"\frac{y-x}{y}=6/15"

"15y-15x=6y"

"y=5\/3x"

differentiate both sides with respect to

t or time.

"dy/dt=(5/3)( dx/dt)=5/3\times5=25/3 ft/sec"

actually, the man's distance from the base doesn't matter since only his speed affects how fast his shadow moves.